# A Comprehensive Survey of Artificial Intelligence Techniques for Talent Analytics

### v2

@@ 2.1.3 Organizational Data
Reporting lines are generally the most representative aspect of an organizational structure, which delineates how authority and responsibility are allocated in an organization. Regarding this point, Sun et al. developed an organization structure-ware convolutional neural network to hierarchically extract compatibility features for measuring person-organization fit and its impact on talent management [381, 383]. Nevertheless, due to privacy concerns, mainstream studies utilize in-firm social networks for human resource management. In general, an in-firm social network can be formed from email or Instant Messaging (IM) records across employees. For example, text-based communication has been used by several machine learning classifiers to identify group mood [213]. Besides, Cao et al. leveraged the lasso regression model to explore team viability using text conversations of online teams [62]. In addition to social networks, researchers have also taken other information into account. For example, Ye et al. utilized both email communication and a high-potential talent list to identify employees with high potential [447]. Along this line, Teng et al. further utilized datasets from three sources for organizational turnover prediction, including profile and turnover, social network, and job levels [386].

### v3

@@ 2.1.3 Organizational Data
Reporting lines are generally the most representative aspect of an organizational structure, as they delineate how authority and responsibility are allocated within an organization. Regarding this point, Sun et al. developed an organization structure-aware convolutional neural network to hierarchically extract compatibility features for measuring person-organization fit and its impact on talent management [25, 26]. Nevertheless, due to privacy concerns, mainstream studies utilize in-firm social networks for human resource management. In general, an in-firm social network can be formed from email or Instant Messaging (IM) records across employees. For example, text-based communication has been used by several machine learning classifiers to identify group mood [63]. Besides, Cao et al. leveraged the lasso regression model to explore team viability using text conversations of online teams [64]. In addition to social networks, researchers have also taken other information into account. For example, Ye et al. utilized both email communication and a high-potential talent list to identify employees with high potential [55]. Along this line, Teng et al. further utilized datasets from three sources for organizational turnover prediction, including profile and turnover, social network, and job levels [65].